# Should Bias Always be Eliminated? A Principled Framework to Use Data Bias for OOD Generation

### v1

@@ Abstract
Most existing methods for adapting models to out-of-distribution (OOD) domains rely on invariant representation learning to eliminate the influence of biased features. However, should bias always be eliminated—and if not, when should it be retained, and how can it be leveraged? To address these questions, we first present a theoretical analysis that explores the conditions under which biased features can be identified and effectively utilized. Building on this theoretical foundation, we introduce a novel framework that strategically leverages bias to complement invariant representations during inference. The framework comprises two key components that leverage bias in both direct and indirect ways: (1) using invariance as guidance to extract predictive ingredients from bias, and (2) exploiting identified bias to estimate the environmental condition and then use it to explore appropriate bias-aware predictors to alleviate environment gaps. We validate our approach through experiments on both synthetic datasets and standard domain generalization benchmarks. Results consistently demonstrate that our method outperforms existing approaches, underscoring its robustness and adaptability.
footnotetext: Correspondence to: Guangyi Chen <guangyichen1994@gmail.com>
@@ 1 Introduction
A widely used example of data bias is the “cow vs. camel classification” problem. As illustrated in Figure 1(a), models trained on imbalanced datasets often rely on the biased background features for prediction. Specifically, cows are typically photographed in grassy fields, while camels appear in desert settings. Consequently, the model may mistakenly learn to associate the background context, rather than the animal itself, with the class label. From a causal perspective, this issue arises due to spurious correlations, where the background acts as a confounding variable influencing both the images and the associated label, as shown in Figure 1(b).
@@
Most existing methods identify data bias as the key challenge in predicting unseen out-of-distribution (OOD) images and thus learn invariant representations to eliminate its influence. For instance, Invariant Risk Minimization (IRM) irm aims to learn a representation that supports a single classifier across all domains. Domain Adversarial Neural Networks (DANN) dann employ adversarial learning to encourage domain-invariant representations. Similarly, Self-Training lee2013pseudo improves generalization by generating pseudo-labels for target domain data based on learned invariant predictions. Some de-bias methods qi2020two leverage causal inference techniques, such as the back-door and front-door rules, to discover true causal relationships from spurious correlations. More discussions on related work can be found in Appendix A.
@@
The core assumption of these methods is that biased features are uninformative, and only features invariant across different domains should be transferred. In this paper, we challenge this assumption and raise a critical question: Should bias always be eliminated? We provide an intuitive motivation for this question in part (c) of Figure 1. As humans, when classifying an image with ambiguous content as either a cow or a camel, we often rely on contextual background cues to make an educated guess. This intuitively suggests that, rather than being entirely detrimental, in some scenarios, the bias derived from background information can sometimes serve as a useful reference in prediction. Mathematically, it indicates that the domain-specific (biased) features and the prediction target are conditionally dependent given the invariant content, which conflicts with the assumption of uninformative bias in invariant representation learning.
Although bias can contain useful information, using it for prediction is challenging and potentially harmful. Recent studies chang2019domain ; wei2018general ; kong2023partial ; erm ; eastwood2024spuriosity have explored the use of biased features, with the most relevant being Stable Feature Boosting (SFB) eastwood2024spuriosity , which adapts conditionally independent biased features using pseudo-labels from stable ones at test time. However, these methods lack theoretical analysis of the conditions under which data bias can be reliably identified and effectively leveraged. Overlooking this can be risky, as bias often contains both “good” components (e.g., context that aids prediction) and “bad” components (e.g., spurious noise that harms prediction), making it essential to carefully disentangle and isolate the useful information.
To address this challenge, we present a theoretical analysis for the identification and utilization conditions of the data bias. Our results show that bias can indeed be leveraged, provided there is sufficient distributional variation across domains/labels and the relationship between the biased features and labels remains unblocked. Notably, even in more general scenarios where domain shifts affect the distributions of both the input data and the label, our analysis demonstrates that biased features can still offer valuable information for improving model performance in the inference stage.
Building on our theoretical analysis, we propose a principled generative framework to identify data bias and effectively leverage it for prediction. Specifically, we employ a Variational Autoencoder (VAE) to learn latent representations and disentangle content features from biased ones in an unsupervised manner. We then leverage the bias for prediction through both direct and indirect strategies. For the direct approach, we extract the beneficial components of the bias by aligning them with invariant predictions. It is achieved by taking the invariant predictions as pseudo-labels to guide the refinement of biased features that filter out harmful bias components, thereby retaining only the useful context for prediction. For the indirect approach, we utilize the biased feature to estimate the underlying environment context and adapt the prediction accordingly. In particular, we decompose the predictor into two components: an invariant predictor that takes content features as input, and a set of domain experts conditioned on biased features. Assuming that the target domain can be represented as a linear combination of source domains, we obtain a bias-aware predictor by the weighted mixture of experts using the inferred bias, enabling better alignment with the target distribution. The main contributions of this paper are summarized as follows:
  • We revisit the role of bias in OOD generalization. Unlike prior work that aims to eliminate bias, we highlight scenarios where bias can be a useful signal for improving prediction.
  • We provide a theoretical analysis to explore conditions under which data bias can be reliably identified and effectively leveraged.
  • We introduce a principled framework to leverage bias in both direct and indirect ways and validate our methods through experiments on both synthetic and real-world datasets.
@@ 2.1 Problem setup and data generation Process
Figure 2: A graphical representation of the data generation process. The content variables ${\mathbf{c}}$ are invariant to environment changes, while the bias variables ${\mathbf{b}}$ vary across environments. Observed data ${\mathbf{x}}$ is generated by $g({\mathbf{c}},{\mathbf{b}})$, with ${\mathbf{c}}$ and ${\mathbf{b}}$ forming the latent variable ${\mathbf{z}}$. The environment $e$ affects ${\mathbf{b}}$ but not ${\mathbf{c}}$, and the target $y$ may also depend on $e$, reflecting complex data-environment interactions.
@@
Notation and problem setup. We formulate the task as a multi-source domain generalization problem, aiming to learn a predictor that remains robust in an unseen target domain using only data from multiple source domains. Let ${\mathbf{x}}_{k}$ denote a high-dimension image observation ${\mathbf{x}}_{k}:=[x_{1},\cdots,x_{n_{x}}]\in{\mathcal{X}}\subset\mathbb{R}^{n_{x}}$, and $y$ denotes the corresponding label. With access to $M$ source domains $\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$, with each domain represented as $({\mathbf{x}}^{\mathcal{S}_{i}},{\mathbf{y}}^{\mathcal{S}_{i}})=({\mathbf{x}}^{\mathcal{S}_{i}}_{k},y^{\mathcal{S}i}_{k})^{m_{i}}_{k=1}$, our goal is to learn a predictor $f({\mathbf{x}})$ that can generalize well to the new domain $\mathcal{T}$ where $\mathcal{T}\nsubseteq\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$.
@@
Data generation Process. To illustrate the origin and impact of bias, we model the data generation process using a latent variable model. Let ${\mathbf{c}}:=[c_{1},\cdots,c_{n_{c}}]\in\mathcal{C}\subset\mathbb{R}^{n_{c}}$ denotes content variables and ${\mathbf{b}}:=[b_{1},\cdots,b_{n_{b}}]\in\mathcal{B}\subset\mathbb{R}^{n_{b}}$ denotes the bias. As shown in Figure 2, the generation process of the observation ${\mathbf{x}}$ is defined as ${\mathbf{x}}:=g({\mathbf{c}},{\mathbf{b}})$, where ${\mathbf{z}}:=[{\mathbf{c}},{\mathbf{b}}]$ with $n_{z}=n_{c}+n_{b}$ dimension denotes the latent variables and $g:[{\mathbf{c}},{\mathbf{b}}]\mapsto{\mathbf{x}}$ denotes the generating function. To distinguish ${\mathbf{c}}$ and ${\mathbf{b}}$, we introduce an environment variable $e$, where the content ${\mathbf{c}}$ remains invariant despite changes of the environment, whereas the bias ${\mathbf{b}}$ varies as the environment changes. Taking the “cow vs. camel classification” problem for illustration, ${\mathbf{c}}$ denotes the characteristics of the object, like the color/shape of cow; while ${\mathbf{b}}$ represents the background information. Note that we consider a more general case where $y$ may also be influenced by the environment, allowing for a more complex relationship modeling between the data and the environment. This assumption is reasonable, as the probability of encountering a cow versus a camel is strongly influenced by the environment (e.g., grasslands vs. deserts) in real-world scenarios. Here we highlight two causal paths between bias and label target, $y\rightarrow{\mathbf{b}}$ and $y\leftarrow e\rightarrow{\mathbf{b}}$, which motivate the direct and indirect ways to leverage the bias for prediction, respectively.
2.2 Theoretical conditions for identifiability and inference
@@
Bias identification. We demonstrate that, with the advancement of representation learning, content and bias variables can be effectively disentangled with an identifiability guarantee. Since disentangling content and bias is fundamental for leveraging biased information, we briefly summarize the mild conditions under which such disentanglement can be achieved. It is important to note that this identifiability result has already been established in previous work and is not claimed as a contribution of this paper. However, despite the identification results, existing methods rely on invariant prediction (including invariant content variable and invariant relation between bias variable and label). Different from them, building on this identifiability framework, we go further by exploring how to effectively utilize biased data to enhance OOD performance.
@@
(Block-wise Identification of ${\mathbf{c}}$ and ${\mathbf{b}}$ kong2023partial .) Assuming that the data generation process follows Figure 2, and the following assumptions hold true:
  • A1 (Smooth and Positive Density): The probability density function of latent variables is smooth and positive, i.e. $P_{\mathbf{z}|\mathbf{e}}>0$.
  • A2 (Conditional Independence): Conditioned on $\mathbf{e}$, each $z_{i}$ is independent of any other $z_{j}$ for $i.j\in\{1,\cdots,n\},i\neq j$, i.e., $\log P_{\mathbf{z}|\mathbf{e},\mathbf{y}}=\sum_{i=1}^{n}q(z_{i},\mathbf{e},\mathbf{y})$, where $q_{i}(z_{i},\mathbf{e},\mathbf{y})$ is the log density of the conditional distribution, i.e., $q_{i}:\log P_{z_{i}|\mathbf{e},\mathbf{y}}$.
@@
  • A3 (Linear Independence): For any $\mathbf{b}\in\mathcal{B}\subseteq\mathbb{R}^{n_{b}}$, there exists $n_{b}+1$ values of $\mathbf{e}$, i.e., $e_{j}$ with $j=0,1,.\cdots,n_{b}$ such that these $n_{b}$ vectors $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})-\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{0})$ with $j=1,\cdots,n_{b}$ are linearly independent, where vector $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}}_{j})$ is defined as follows:
    $\bm{{\mathbf{w}}}({\mathbf{b}},{\mathbf{e}})=\left(\frac{\partial q_{1}({\mathbf{b}}_{1},{\mathbf{e}})}{\partial{\mathbf{b}}_{1}},\cdots\frac{\partial q_{n_{b}}({\mathbf{b}}_{n_{b}},{\mathbf{e}})}{\partial{\mathbf{b}}_{n_{b}}}\right),$ (1)

### v2

# Should Bias be Eliminated? A General Framework to Use Bias for OOD Generalization

@@ Abstract
Most approaches to out-of-distribution (OOD) generalization learn domain-invariant representations by discarding contextual bias. In this paper, we raise a critical question: Should bias be eliminated? If not, is there a general way to leverage bias for better OOD generalization? To answer these questions, we first provide a theoretical analysis that characterizes the circumstances in which biased features contribute positively. Although theoretical results show that bias may sometimes play a positive role, leveraging it effectively is non-trivial, since its harmful and beneficial components are often entangled. Recent advances have sought to refine the prediction of bias by presuming reliable predictions from invariant features. However, such assumptions may be too strong in the real world, especially when the target also shifts from training to testing domains. Motivated by this challenge, we introduce a framework to leverage bias in a more general scenario. Specifically, we employ a generative model to capture the data generation process and identify the underlying bias factors, which are then used to construct a bias-aware predictor. Since the bias-aware predictor may shift across environments, we first estimate the environment state to train predictors under different environments, combining them as a mixture of domain experts for the final prediction. Then, we build a general invariant predictor, which can be invariant under label shift to guide the adaptation of the bias-aware predictor. Evaluations on synthetic data and standard domain generalization benchmarks demonstrate that our method consistently outperforms both invariance only baselines, recent bias utilization approaches and advanced baselines, yielding improved robustness and adaptability.
Machine Learning, ICML
@@ 1 Introduction
A widely used example of data bias is the “cow vs. camel classification” problem. As illustrated in Figure 1(a), models trained on imbalanced datasets may rely on biased background features for prediction. Specifically, cows are typically photographed in grassy fields, while camels appear in desert settings. Consequently, the model may mistakenly learn to associate the background context, rather than the animal itself, with the class label. From a causal perspective, this issue arises from spurious correlations, where the background acts as a confounding variable influencing both the images and the associated label, as shown in Figure 1(b).
@@
Most OOD generalization methods, such as IRM (Arjovsky et al., 2019), DANN (Ganin et al., 2016), and Self-Training (Lee and others, 2013), regard data bias as the primary obstacle and therefore aim to learn domain-invariant representations (see Appendix A for further discussion). The core assumption of these methods is that biased features are uninformative and that only invariant features should be transferred. As illustrated in Figure 1(c), when an image is ambiguous (e.g., cow vs. camel), humans often rely on background context to make an educated guess. This intuitively suggests that, rather than being entirely detrimental, in some scenarios, the bias derived from background information can serve as a useful reference in the prediction process.
To analyze the effect of bias, we provide a theoretical framework that characterizes when data bias is identifiable and beneficial for prediction. Our results show that, in general, biased features can be exploited when they can be disentangled from invariant content, and the dependence between biased features and labels is not screened off by invariant content. It indicates that we can disentangle the bias component and study its predictive contribution, enabling bias features to provide complementary information that can be effectively exploited at inference.
Although bias contains useful information, using it for prediction is challenging. The most relevant prior work that leverages bias for OOD generalization is Stable Feature Boosting (SFB)  (Eastwood et al., 2023). SFB splits the representation into invariant and biased components, using the predictions based on the invariant features as pseudo-labels for test-time adaptation of the biased component. However, this method relies on the assumption that invariant features alone can yield accurate and robust predictions. In particular, it presumes that (1) invariant features contain sufficient information to support prediction, and (2) the prediction target itself remains invariant, i.e., no label shift (Lipton et al., 2018) occurs. In real-world applications, the invariant content may be under-learned due to limited domain heterogeneity, insufficient supervision, or an intrinsic lack of information. Meanwhile, the prediction using the invariant feature may shift in the testing domain, as shown in the common label shift setting (Li et al., 2023b; Wu et al., 2021; Koh et al., 2021).
To leverage bias in more general and realistic settings, we adopt a generative model that factorizes representations into invariant content and contextual bias in an identifiable manner. Building on the identified bias features, we uncover two pathways through which bias can contribute to prediction. Firstly, we use the bias features to estimate environmental states and design an environment routing mechanism that leverages these states to gate a mixture of predictors, each conditioned on a specific environment. This environment estimation enables the model to exploit bias as a useful signal, particularly in cases where the invariant content alone is insufficient for accurate prediction. Secondly, we extend the bias correction in SFB (Eastwood et al., 2023) with an adaptive label prior. We compose predictors using bias and content together by Bayes’ Rule into two parts: bias predictor and invariant predictor which consists of content predictor and label prior. In addition to correcting the bias predictor using content predictions as pseudo-labels with a fixed label prior, we further learn the label prior in the training stage. This modification can effectively address the problem when both label shift and covariate shift co-exist. The contributions of this paper are summarized as follows:
  • We revisit the role of bias in OOD generalization, highlighting circumstances in which bias offers a constructive signal rather than a nuisance to be eliminated.
  • We provide a theoretical analysis that specifies identification and utility conditions for data bias, including cases with concurrent shifts in input and label distributions.
  • We introduce a general bias-aware framework that makes predictions with an environment routing mechanism and conducts adaptive bias correction.
  • We evaluate the proposed method on synthetic and real-world datasets and show that it consistently outperforms baseline approaches across both settings.
@@ 2.1 Problem Setup and Data Generation Process
Notation and problem setup. We formulate the task as a multi-source domain generalization problem, aiming to learn a predictor that remains robust in an unseen target domain using only data from multiple source domains. Let ${\mathbf{x}}_{k}$ denote a high-dimensional image observation ${\mathbf{x}}_{k}:=[x_{1},\cdots,x_{n_{x}}]\in{\mathcal{X}}\subset\mathbb{R}^{n_{x}}$, and $y_{k}$ denote the corresponding label. With access to $M$ source domains $\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$, with each domain represented as $({\mathbf{x}}^{\mathcal{S}_{i}},{\mathbf{y}}^{\mathcal{S}_{i}})=({\mathbf{x}}^{\mathcal{S}_{i}}_{k},y^{\mathcal{S}_{i}}_{k})^{m_{i}}_{k=1}$, our goal is to learn a predictor $f({\mathbf{x}})$ that can generalize well to the new domain $\mathcal{T}$ where $\mathcal{T}\notin\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{M}\}$.
Figure 2: A graphical representation of the data generation process. The content variables $\mathbf{c}$} are invariant to environment changes once label $y$ is given, while the bias variables $\mathbf{b}$} vary across environments. Observed data $\mathbf{x}$} is generated by $g({\mathbf{c}},{\mathbf{b}})$, with $\mathbf{c}$} and $\mathbf{b}$} forming the latent variable $\mathbf{z}$}. The environment $e$ affects $\mathbf{b}$} but not $\mathbf{c}$}, and the target $y$ may also depend on $e$, reflecting complex data-environment interactions.
@@
Data generation process. To illustrate the origin and impact of bias, we model the data generation process using a latent variable model. Let ${\mathbf{c}}:=[c_{1},\cdots,c_{n_{c}}]\in\mathcal{C}\subset\mathbb{R}^{n_{c}}$ denote content variables and ${\mathbf{b}}:=[b_{1},\cdots,b_{n_{b}}]\in\mathcal{B}\subset\mathbb{R}^{n_{b}}$ denote the bias. As shown in Figure 2, the generation process of the observation $\mathbf{x}$} is defined as ${\mathbf{x}}:=g({\mathbf{c}},{\mathbf{b}})$, where ${\mathbf{z}}:=[{\mathbf{c}},{\mathbf{b}}]$ with $n_{z}=n_{c}+n_{b}$ dimensions denotes the latent variables and $g:[{\mathbf{c}},{\mathbf{b}}]\mapsto{\mathbf{x}}$ denotes the generating function. To distinguish $\mathbf{c}$} and $\mathbf{b}$}, we introduce an environment variable $e$, where the content $\mathbf{c}$} remains invariant despite changes in the environment once $y$ is given, whereas the bias $\mathbf{b}$} varies as the environment changes. Taking the “cow vs. camel classification” problem for illustration, $\mathbf{c}$} denotes the characteristics of the object, like the color/shape of the animal, while $\mathbf{b}$} represents the background information.
Note that we consider a more general case where $y$ may also be influenced by the environment, allowing for a more complex relationship modeling between the data and the environment. This assumption is reasonable, as the probability of encountering a cow versus a camel is strongly influenced by the environment (e.g., grasslands vs. deserts) in real-world scenarios. Here we highlight two causal paths between bias and label target, $y\leftarrow e\rightarrow{\mathbf{b}}$ and $y\rightarrow{\mathbf{b}}$, which motivate two ways to leverage the bias for prediction. First, we can estimate the environment state from $\mathbf{b}$} and use such information to adjust the prediction. Second, we can leverage content to guide the direct prediction from $\mathbf{b}$} to $y$.
2.2 Theoretical Analysis For Identifiability and Usage
Bias and Content identification.
@@ 2.2 Theoretical Analysis For Identifiability and Usage
We demonstrate that content and bias variables can be effectively disentangled with an identifiability guarantee. We extend the theoretical results of the previous causal representation learning methods to align with our formulation, with the detailed modification, proof and discussions that can be found in Sections B.1 and B.2. Such block-wise identification enforces invertible mappings between learned and ground-truth factors: $\hat{\mathbf{c}}=f({\mathbf{c}})$, $\hat{\mathbf{b}}=g({\mathbf{b}})$, with $f,g$ being invertible. It is important to note that this identifiability result has already been established in previous work and is not claimed as a contribution of this paper. However, despite the identification results, existing methods still rely on invariant prediction (including invariant content variable and invariant relation between bias variable and label). Unlikely, building on this identifiability framework, we go further by exploring how to effectively utilize biased data to enhance OOD performance.
@@
(Block-wise Identification of $\mathbf{c}$} and $\mathbf{b}$} (Kong et al., 2022)). Assuming that the data generation process follows Figure 2, and the following assumptions hold true:
  • A1 (Smooth and positive density): $p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})$ is smooth with strictly positive density on its support.
  • A2 (Conditional independence and content invariance): Given $(\mathbf{e},\mathbf{y})$, the coordinates of $\mathbf{z}$ are independent, $\log p_{\mathbf{z}\mid\mathbf{e},\mathbf{y}}(\mathbf{z}\mid\mathbf{e},\mathbf{y})=\sum_{i=1}^{n_{z}}q_{i}(z_{i};\mathbf{e},\mathbf{y}),$
@@
  • A3 (Linear independence): For any $\mathbf{b}\in\mathcal{B}\subseteq\mathbb{R}^{n_{b}}$ (and the relevant $\mathbf{y}$), there exist $n_{b}+1$ environments $e_{0},e_{1},\ldots,e_{n_{b}}$ such that the $n_{b}$ vectors $\bm{w}(\mathbf{b},e_{j},\mathbf{y})-\bm{w}(\mathbf{b},e_{0},\mathbf{y}),\quad j=1,\ldots,n_{b},$ are linearly independent, where $\bm{w}(\mathbf{b},e_{i},\mathbf{y}):=\Big(\tfrac{\partial q_{1}(b_{1};e_{i},\mathbf{y})}{\partial b_{1}},\ldots,\tfrac{\partial q_{n_{b}}(b_{n_{b}};e_{i},\mathbf{y})}{\partial b_{n_{b}}}\Big)^{\top}.$